OpenStudy (anonymous):

Differentiate y= (cos^-1 3x)^3

OpenStudy (anonymous):

Gotta use the chain rule.

OpenStudy (anonymous):

\[\frac{d}{dx}\left[(cos^{-1}(3x))^3\right]\]\[=3(cos^{-1}(3x))^2\cdot\frac{d}{dx}\left[cos^{-1}(3x)\right]\]\[=3(cos^{-1}(3x))^2\cdot\frac{-1}{\sqrt{1-(3x)^2}}\cdot\frac{d}{dx}[3x]\] \[=-\frac{9(cos^{-1}(3x))^2}{\sqrt{1-9x^2}}\]